Good day madam/sir.
can you please help me out on my assignment i need this tomorrow and probably i can't understand this. i have so many questions here but can you help me with this only? thank you.
Some of Aron's friends are planning to buy him a birthday gift worth 270 pesos, dividing the cost equally among themselves. six more of his friends decided to share in the expenses and so each one's share is decreased by 12 pesos. how many friends were originally part of the plan.

Answer by mananth(16946)

let there be x friends initially
6 joined them
so total = x+6
Initially each had to pay 270/x pesos
Later their contribution became 270/(x+6)
270/x - 270/(x+6) = 12
270(x+6)-270x = 12x(x+6)
270x+1620-270x=12x^2+72x
12x^2+72x-1620=0
/12
x^2+6x-135=0
x^2+15x-9x-135=0
x(x+15)-9(x+15)=0
(x-9)(x+15)=0
x=9 OR -15
there were 9 friends initially